OpenStudy (like1234):

could someone tell me the answer for this problem (34 + 14) ÷ [42 – (20 + 16)]

OpenStudy (cj49):

34+14=48 42-(36)=6 48/6=8

OpenStudy (like1234):

thanks

OpenStudy (cj49):

np